Calcium reacts with water to produce hydrogen gas and aqueous calcium hydroxide. (a) Write the balanced equation for the reaction. (b) How many mL of water (d= 1.00 g/mL) are required to produce 7.00 L of dry hydrogen gas at 1.05 atm and 32°C? 


1
Expert's answer
2021-12-13T17:07:02-0500

H2O + Ca → H2 + CaO

Molar mass of H2= 2.016

7000/2.016=3472.22

Molar mass H2O=18.012

=3472.22×18.012=62541.67ml